What potential difference is required to accelerate a proton from rest to a speed of 5.72E7 m/s?




1
Expert's answer
2021-05-30T13:29:25-0400
"PE=KE,""q\Delta V=\dfrac{1}{2}mv^2,""\Delta V=\dfrac{mv^2}{2q}""\Delta V=\dfrac{1.67\cdot10^{-27}\ kg\cdot(5.72\cdot10^7\ \dfrac{m}{s})^2}{2\cdot1.6\cdot10^{-19}\ C}=1.7\cdot10^7\ V."